OpenStudy (anonymous):

Solve the equation, using the domain of x = {-1, 0, 2}. 2x + 3y = 7

OpenStudy (anonymous):

hey vi3tbloodxx values of x are in sets not complete range of values so how can we get range of values for y???

OpenStudy (anonymous):

none of those are any of the oppions

OpenStudy (anonymous):

it actually x = {-1, 0, 2} 2x+3y=7 2(-1)+3y=7 -2+3y=7 3y=9 y=3 (-1,3) 2x+3y=7 2(0)+3y=7 3y=7 y=7/3 (0,7/3) 2x+3y=7 2(2)+3y=7 4+3y=7 3y=3 y=1 (2,1)

OpenStudy (anonymous):

Oh my bad soz, got a lil sleepy saw it wrong.

OpenStudy (anonymous):

it fine dont worry about